OpenStudy (anonymous):

The retina of the human eye can detect light when radiant energy incident on it is at least 4.0x10^-17

OpenStudy (anonymous):

Use E = hf f = speed of light/wavelength = 3*10^8 / 600*10^-9 h = 6.62 * 10^-34 Therefore, E = hf = 3.31*10^-19 J The min energy is 4*10^-17 J Therefore, 4.0 x10^-17 / 3.31*10^-19 J = 120 photons.